I submitted a couple of days ago but didn't get a response. Can anyone help me?
10x^2 + x = 21
use the discriminant to determine the nature of the roots for equation. do not solve equation.
thank you...

Answer by checkley71(8403)

10X^2+X=21
10^2+X-21=0
(5X-7)(2X+3)=0
5X-7=0
5X=7
X=7/5
AND
2X+3=0
2X=-3
X=-3/2